Events A and B are independent events.
Find the indicated Probability.
P (A) = 0.4
P (B) = 0.7
P (A and B) = _______

Answer by ikleyn(52943)

.

Since the events A and B are independent,


    P(A and B) = P(A) * P(B) = 0.4 * 0.7 = 0.28.      ANSWER

Solved.